Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
pinto, ruggi, gherardini - 19:22 Thursday 20 June 2024 (64557) Print this report
test of DARM and DIFFp TY control filters

Today's 1 hr shift was meant to resume the leftover activities concerning the LF optimization. In details we wanted to try:

1) boost control filter for DARM (improve the LF loop accuracy of a factor 2 up to 5 Hz);

2) DIFFp TY rolloff to reduce the control noise reintroduction starting from 15-20 Hz, since from the last ASC projections we noticed an excess of coupling of such loop with Hrec.

At the beginning of the shift (17.04) we first tried the new control filter for DARM. We unlocked immediately after the start of the ramp of the filter swap. We used a 100 second ramp since in the past when we tried a new DARM filter with a shorter ramp (i.e. 5 sec), we unlocked immediately probably due to a too quick discharge/charge of the integrator of the filter. By using a longer ramp we avoided this issue. However this time it didn't work. By looking at the data and at the filter gain the reason of the unlock is still not clear. Data will be better analyzed.

Profiting of the unlock we loaded the new DIFFp TY control filter and we tried it from the beginning of the loop engagement (CARM null 3f). The filter engagement worked fine, however we didn't have the chance to see its effect in LN3 since the ITF kept unlocking during the ACQUIRE LN3 step. The reason of the unlock is not entirely clear, we only noticed a weird behaviour of BS TY loop right before the unlock (jumps on the error signal) followed by a flash on B1p DC, see Fig.1 reporting the unlock of 16.23.53 for reference. We will look carefully at the data.

We reverted the filter of DIFFp TY to the nominal one and left the ITF locking for the next activities.


Images attached to this report
Comments to this report:
pinto - 11:17 Friday 21 June 2024 (64561) Print this report

Yesterday it has been loaded a wrong filter for DIFFpTY with much more roll-off, but less stability in terms of gain and phase margins (fig.1, blue current filter , red new wrong filter).

During the transition to LN3, the gain servo increased the gain loop of about a factor 1.4, to adjust some loos of gain. Even if we had less gain margin (and a factor +1.4 of gain would have possibly brought the loop to the upper limit of phase crossing ), this doesn't explain the unlocks, also because in the data we didn't experience visible loop oscillation due to high loop gain (Fig2 and Fig 3 the two unlocks at ACQUIRE LN3 steps).


Images attached to this comment
Search Help